OpenStudy (anonymous):

What is the value of the range of the function f(x) = x² + 2 for the domain value 1/4 ?

OpenStudy (dumbcow):

plug in 1/4 for x --> (1/4)^2 +2

OpenStudy (anonymous):

Substitute 1/4 (1/4)^2 + 2 = 2 1/16

OpenStudy (anonymous):

A) y = 2 B) y = 2/4 C) y = 21 /8 D) y = 21/16

OpenStudy (anonymous):

Are you sure about the options?

OpenStudy (anonymous):

D 2 1/16

OpenStudy (anonymous):

meverett04, good guess however D looks like \(\large\frac {21}{16} \) rather than \( \large2\frac1{16}\) (to me).